Riv. Mat. Univ. Parma, Vol. n (202x), 000-000

FirstName1 LastName1, FirstName2 LastName2 and FirstName3 LastName3

Title of the article

Abstract. This is the text of the abstract ….

Keywords. Word1, word2, word3, wordn.

Mathematics Subject Classification: xxaayy, xxaayy, xxaayy.

1 - Title of section Text of the section …. 1.1 - Title of subsection Text of the subsection ….

1.1.1 - Title of subsubsection

Text of the subsubsection ….

(1.1)  lim_{r→∞} ∫_{r}^{∞} v^{s}f(v) dv = 0.

D e f i n i t i o n 1.1. Here is the text of the definition. Use typeface emph for the name of the term defined.

P r o p o s i t i o n 1.2. Here is the text of the proposition ….

This research was partially supported by ............

P r o o f. Here is the proof of the proposition …. □

A c k n o w l e d g m e n t s. The authors acknowledge support by the ….

References

[1] N. Author1, N. Author2 and N. Author3, Title of the article, Journal abbreviation vol.number (year), pp-pp.

[2] N. Author1 and N. Author2, Title of the book, Publishing House, Town, Year.

First Author University of Xxxxxxxxx, Address Address City, Zip code, Country e-mail: xxxxxxxxx@yyyyyy.zz

Second Author University of Xxxxxxxxx, Address Address City, Zip code, Country e-mail: xxxxxxxxx@yyyyyy.zz

Third Author University of Xxxxxxxxx, Address Address City, Zip code, Country e-mail: xxxxxxxxx@yyyyyy.zz